OpenStudy (anonymous):

1/2 (4n-6) = 11 Solve for n

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Yes @mathstudent55

Directrix (directrix):

@mysticalfire 1/2 (4n-6) = 11 Multiply both sides of the equation by 2. 2 * 1/2 (4n-6) = 2* 11 [2 * 1/2] (4n-6) = 2 * 11 What step comes next?